# LLaMA-Berry: Pairwise Optimization for O1-like Olympiad-Level Mathematical Reasoning

### v1

@@ 2.4 Enhanced Borda Count Method
Figure 3: The Pipeline of Enhanced Borda Count (EBC) Method
@@
Although PPRM allows us to directly compare the quality of two answers, we still need to convert these local partial order relations into a cohesive global ranking to gain a comprehensive evaluation for the answers. This conversion process can be formalized as the global optimal ranking aggregation (GORA) problem. To address this problem, we propose the Enhanced Borda Count (EBC) method based on a transitivity assumption of mathematical problem-solving solutions (detailed in Figure 3), which integrates the naive Borda Count Algorithm with a transitive closure of partial order relations calculated by the Floyd-Warshall (Cormen et al., 2009) Algorithm. For formalized proof, please refer to Appendix E.
@@
Local Preference Calculation: First, as shown in Figure 3 (1), the PPRM generates a win-loss matrix $M\in\mathbb{R}^{n\times n}$ for all $n$ problem-solving solutions, where $M[i,j]=1$ indicates that solution $a_{i}$ is superior to solution $a_{j}$, and $M[i,j]=0$ otherwise. This process can be represented as:
$M[i,j]=\begin{cases}1,&\text{if }P(a_{i}\succ a_{j})\geq 0.5\\ 0,&\text{if }P(a_{i}\succ a_{j})<0.5\end{cases}$ (6)
@@
This matrix can be viewed as an adjacency matrix of a directed graph $G=(V,E)$, where each solution $a_{i}$ corresponds to a vertex $v_{i}$, and an edge $e=(v_{i},v_{j})$ exists if $M[i,j]=1$, indicating that solution $a_{i}$ is preferred over $a_{j}$.
@@
Transitive Closure: As shown in Figure 3 (2), to simplify the problem, we adopt the assumption of transitivity for mathematical solutions, that is, if $v_{i}\succ v_{k}$ and $v_{k}\succ v_{j}$, then $v_{i}\succ v_{j}$. Under this assumption, the transitive closure $C$ of a preference matrix can be computed. Specifically, if $M[i,k]=1$ and $M[k,j]=1$, then $M[i,j]=1$. This is achieved by updating the transitive closure matrix $C$ using Floyd-Warshall algorithm with dynamic programming rule:
$C[i,j]=\bigvee_{k=1}^{n}(C[i,k]\wedge C[k,j])$ (7)
where $\bigvee$ denotes the logical OR operator and $\wedge$ denotes the logical AND operator.
@@
Borda Count-Based Global Ranking: Next, based on the transitive closure matrix $C$, we apply the Borda Count method for global ranking. The Borda Count determines the ranking of each node by calculating its out-degree, which corresponds to the number of nodes it defeats. For each node $v_{i}$, the Borda Count $\text{Borda}(v_{i})$ is defined as $\sum_{j=1}^{n}C[i,j]$, like the ranked node list in Figure 3 (3).
@@
Nodes with higher Borda counts are ranked higher. However, in practice, cyclic preferences can cause efficiency issues with the naive Borda Count method. To further refine the ranking, as shown in Figure 3 (4), a re-ranking stage is introduced, where the logits generated by the PPRM are used for soft comparison among nodes with equal Borda counts. Specifically, for two nodes $v_{i}$ and $v_{j}$ with equal Borda counts, the soft comparison rule can be denoted as $v_{i}\succ v_{j}\iff P(a_{i}\succ a_{j})>P(a_{j}\succ a_{i})$. This process ensures that even in the presence of cycles or local ambiguities, the ranking remains consistent and reasonable.
@@
Global Quantile Score of Solutions: Finally, as shown in Figure 3 (5), the ranking is converted into a global quantile quality score $Q_{g}$ for each solution $v$ is $Q_{g}[v]=1-\frac{\text{rank}(v)-1}{|V|-1}$, where $\text{rank}(v)$ is the position of $v$ in the ranking based on Borda counts, and $|V|$ is the total number of nodes. To reflect local advantages in the search tree structure, the local win rate $Q_{l}(x)$ for a node $v$ is calculated with its children nodes $\text{Children}[v]$ as follows:
@@
The final score $Q(x)$ for a solution is a weighted sum of local win rate $Q_{l}(x)$ and global ranking quantification $Q_{g}$, computed as:
$Q(x)=\alpha\cdot Q_{l}(x)+(1-\alpha)\cdot Q_{g}(x)$ (9)
where $\alpha$ controls the balance between local win probabilities and global ranking. The softened re-ranking process ensures that even in cyclic preferences, the local win probabilities and global ranking results are consistent, ultimately producing a reasonable and reliable ranking.
@@ 3.1 Experiment Settings
Settings: To better evaluate the effectiveness of our approach, we select LLaMA-3.1-8B-Instruct model (Meta, 2024b) as the base model for SR-MCTS search, without any additional training. We train a Gemma2-2B-Instruct model (Google, 2024) as PPRM to provide reward signals during the search process. To ensure robust and efficient inference, we contribute the Berry-Tree inference framework to the community, which supports advanced features such as fault tolerance, checkpoint recovery, multi-query concurrency, and automatic multi-server load balancing (hyper-parameter settings detailed in Appendix A).
@@
Grading: We assess algorithm-generated answers using the correctness evaluation method from PRM800K (Lightman et al., 2023), focusing on format adherence and content accuracy. The model receives a prompt detailing the expected answer format. We score answers as consistent if they exactly match the ground truth, closely approximate it numerically, or are equivalent in symbolic form. Then, to provide a comprehensive and robust evaluation metric, we adopt major@k and pass@k based on our reward-ranked sorting which is denoted as rm@k, as the evaluation metrics. Details on the evaluation method and metrics are further outlined in the Appendix B.
3.2 Benchmark
Table 1: Benchmark Comparison of Different Math Models
DeepSeekMath-7B-RL (Shao et al., 2024) 88.2 52.4 43.6 19.0 37.5 64.8
DeepSeek-Coder-V2-Lite-Instruct (Zhu et al., 2024) 87.6 61.0 56.1 26.4 39.8 68.6
Internlm2-math-plus-20b (Ying et al., 2024) 87.9 56.5 51.9 23.1 37.5 63.5
Internlm2-math-plus-mixtral8x7b (Ying et al., 2024) 92.1 59.4 49.6 25.0 37.5 71.9
Mathstral-7B-v0.1 (Mistral AI, 2024) 84.9 56.6 46.0 21.5 33.7 64.0
Qwen2-7B-Instruct (Yang et al., 2024) 85.7 52.9 36.4 21.3 24.5 68.2
Qwen2-Math-7B-Instruct (Yang et al., 2024) 89.9 75.1 62.1 38.2 45.9 63.8
Qwen2-72B-Instruct (Yang et al., 2024) 93.2 69.0 58.7 33.2 43.2 84.4
Qwen2-Math-72B-Instruct (Yang et al., 2024) 96.7 84.0 68.3 43.0 47.9 79.9
LLaMA 3.1-8B-Instruct (Meta, 2024b) + SR-MCTS (Ours) 89.8@major8 54.8@major8 36.4@major8 24.8@major8 36.4@major8 68.3@major8
94.9@rm8 69.4@rm8 61.6@rm8 47.2@rm8 63.7@rm8 82.9@rm8
96.1@rm16 75.3@rm16 68.6@rm16 55.1@rm16 68.9@rm16 88.3@rm16
@@ 3.2 Benchmark
General Mathematical Reasoning Benchmarks We summarize the results on general mathematical reasoning benchmarks in Table 1, which indicates that our method significantly boosts the base model’s performance. The results also illustrate that the improvements are consistent across various difficulty levels. Specifically, the capability of Meta-Llama-3.1-8B-Instruct has improved by more than 35% on 4 benchmarks at rm@16. Qwen2-Math-72B-Instruct has demonstrated the strongest mathematical capability among the competing methods, while our LLaMA-Berry exceeds it on four benchmarks. In particular, LLaMA-Berry reaches 55.1% on OlympiadBench and 68.9% on College Math, surpassing the 70B model by 11.9% and 21%, respectively.
Additionally, we observe that majority voting results are often lower than rm@k under the same number of simulations. For benchmarks like GSM8K and MATH, The performance of majority voting is not significantly inferior to that of rm@k. However,for those more challenging (e.g., GaoKao 2023 En, OlympiadBench and College Math shown in Table 1), the performance gap between their result become wider. We suspect the reason is that, on benchmarks with lower question difficulty, the performance of LLM primarily relies on its inherent reasoning capabilities, whereas on more complex benchmarks , it’s performance is largely dependent on its Self-Refine abilities. As shown in Table 2, in both the AMC2023 and GPQA Diamond benchmarks, we again observe this phenomenon, which becomes increasingly evident as benchmark difficulty scales when comparing MATH and other benchmarks. This indicates the potential of our method on solving more complex reasoning problems
A comparison between MATH and OlympiadBench shows that the latter is much more difficult and challenging, with less dataset leakage and higher inference costs, making Self-Refine harder to apply. The increased variance in answer sequences means the model does not hit the diminishing returns ceiling of Self-Refine as quickly as it does with GSM8K.
@@
Table 2: Performance Comparison across Multiple Olympiad Benchmarks
@@
MATH AIME2024 AMC2023 Math Odyssey GPQA Diamond TheoremQA DeepMind Mathematics AIME1983-2023
Claude 3 Opus 60.1 6.7 42.0 (5-shot) 40.6 50.4
GPT 4 Turbo 73.4 3.3 47.0 38.8 48.4
GPT 4o 60.3 13.4 56.1
OpenAI o1 Preview 85.5 56.7 78.3
OpenAI o1 94.8 83.3 78.0
Gemini 1.5 Pro 67.7 6.7 45.0 46.2
Gemini Math-Specialized 1.5 Pro 80.6 23.3 55.8
Meta-LLaMA-3.1-8B-Instruct 47.2 6.7 15.7 41.7 30.4 25.7 49.3 15.5
Qwen2-Math-7B-Instruct 75.1 13.3 62.5
NuminaMath-72B CoT 66.7 3.3 52.5
Qwen2-Math-72B-Instruct 84.0 20.0 60.0
LLaMA-3.1-8B-Instruct+ SR-MCTS 54.8@major8 13.3@major8 22.9@major8 44.2@major8 39.4@major8 29.7@major8 56.4@major8 18.0@major8
69.4@rm8 16.7@rm8 48.2@rm8 60.4@rm8 77.3@rm8 39.9@rm8 60.1@rm8 28.6@rm8
75.3@rm16 26.7@rm16 54.2@rm16 65.0@rm16 92.4@rm16 45.5@rm16 63.3@rm16 34.8@rm16
Cutting-edge Mathematical Olympiad Benchmarks: We evaluate the performance of our model on the cutting-edge mathematical olympiad benchmarks, as shown in Table 2. The results demonstrate that LLaMA-Berry remains highly competitive on Olympic-level problems, demonstrating its capability in complex reasoning. Notably, on the most challenging AIME2024 benchmark, the success rate increases from 2/30 (LLaMA-3.1-8B-Instruct) to 8/30 (ours), which not only surpasses typical open-source models but also exceeds all commercial closed-source solutions, except the OpenAI o1 series. On the GPQA Diamond benchmark, which includes advanced graduate-level questions across physics, biology, and chemistry—representing the highest reasoning difficulty. Our algorithm achieves accuracy comparable to GPT-4 Turbo on this benchmark.
Our approach demonstrates remarkable performance not only on mathematical reasoning problems but also across a variety of science and engineering problems. For instance, it excels in benchmarks such as GPQA diamond in Table 1 and MMLU STEM in Table 2. This highlights the generalizability and robustness of our method, which can effectively tackle a wide range of challenges in technical domains, thereby showcasing its potential for broader applications in both research and practical scenarios.
Table 3: Performance of Different Tree-Based Methods in LLaMA-3-8B-Instruct on GSM8K, GSMHARD, and MATH500 Benchmarks
         Method Benchmark         GSM8K (%)         GSMHARD (%)         MATH500 (%)
@@
        Self-Consistency with 8 Samplings         78.4         28.5         30.0
        Self-Consistency with 64 Samplings         83.2         30.3         33.0
        Self-Consistency with 128 Samplings         84.7         31.2         33.8
        Tree of Thoughts         69.1         19.6         13.6
        Reasoning via Planning         80.6         29.6         18.8
        rStar with 32 Rollouts         88.7@major32         33.4@major32         38.3@major32
        rStar with mutual reasoning and 32 Rollouts         91.7@rm32         37.5@rm32         42.9@rm32
        SR-MCTS with 8 Rollouts         86.4@major8         30.2@major8         35.2@major8
        94.1@rm8         37.1@rm8         56.4@rm8
        SR-MCTS with 16 Rollouts         88.1@major16         31.5@major16         39.6@major16
        96.4@rm16         41.1@rm16         63.8@rm16
Comparison with Other Tree-Based or CoT Methods: We also compare our algorithm with other tree-based reasoning methods and CoT on GSM8K, GSMHard (Gao et al., 2022), and MATH500 (Lightman et al., 2023), a representative and highly challenging 10% subset of MATH. The results are depicted in Table 3. Shown in Table 3, as the benchmark difficulty increases (from GSM8K to GSMHard), performances of the RAP and ToT methods tend to degrade relative to simpler methods like Few-shot CoT and One-turn Self-Refine. We suspect the reason could be the weak self-evaluation capability of LLMs, which may guide reasoning step to wrong side. Moreover, tree search-based methods can incur a more computational overheads compared to these simpler methods. In contrast, rStar and our method maintain a positive trend in output performance, highlighting the superior energy efficiency of both approaches.
By analyzing the result from rStar, Self-consistency, and our algorithm, it becomes evident that our method achieves the highest efficiency in rollouts. Specifically, when observing majority voting metrics, our method achieves an accuracy of 86.4% which is among the same level on GSM8K as others, with only 1/4 of the exploration cost of rStar and 1/16 of Self-consistency. This provides compelling validation of the efficacy of the global evaluation mechanism within PPRM, as well as the aggressive exploration fostered by the dynamic pruning strategy.

### v2

@@ 2.4 Enhanced Borda Count Method
Although PPRM allows us to directly compare the quality of two solutions, we still need to convert these local preferences into a cohesive global ranking to gain a comprehensive evaluation for the answers. This conversion process can be formalized as the global optimal ranking aggregation (GORA) problem related to Learning to Rank (LtR) methods. Further, we propose the Enhanced Borda Count (EBC) method based on the transitivity assumption of mathematical problem solutions, which integrates the naive Borda Count algorithm with a transitive closure of preferences calculated by the Floyd-Warshall (Warshall, 1962) algorithm. For formalized discussion, please refer to Appendix H.
@@
Local preference calculation. First, the PPRM generates a preference matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ for all $n$ problem solutions, where $\mathbf{M}_{i,j}=1$ indicates that solution $a_{i}$ is superior to solution $a_{j}$, and $\mathbf{M}_{i,j}=0$ otherwise. This process can be represented as:
$\mathbf{M}_{i,j}=\begin{cases}1,&\text{if }P(a_{i}\succ a_{j})\geq 0.5\\ 0,&\text{if }P(a_{i}\succ a_{j})<0.5\end{cases}$ (6)
@@
As shown in Figure 2, this matrix can be viewed as an adjacency matrix of a directed graph $\mathbf{G}=(\mathbf{V},\mathbf{E})$, where each solution $a_{i}$ corresponds to a vertex $v_{i}$, and an edge $e=(v_{i},v_{j})$ exists if $\mathbf{M}_{i,j}=1$, indicating that solution $a_{i}$ is preferred over $a_{j}$.
@@
Transitive closure. To simplify the problem, we adopt the assumption of transitivity for mathematical solutions, that is, if $v_{i}\succ v_{k}$ and $v_{k}\succ v_{j}$, then $v_{i}\succ v_{j}$. Under this assumption, the transitive closure $\mathbf{C}$ of a preference matrix can be computed through Floyd-Warshall algorithm, e.g., if $\mathbf{M}_{i,k}=1$ and $\mathbf{M}_{k,j}=1$, then $\mathbf{M}_{i,j}=1$.
@@
Borda count-based global ranking. Next, based on the transitive closure matrix $\mathbf{C}$, we apply the Enhanced Borda Count method for global ranking. The Enhanced Borda Count determines the ranking of each node by calculating its out-degree, which corresponds to the number of nodes it defeats. For each node $v_{i}$, the $\text{Borda}(v_{i})$ is defined as $\sum_{j=1}^{n}\mathbf{C}_{i,j}$, like the ranked node listed in Figure 2.
@@
Nodes with higher Borda counts are ranked higher. However, in practice, cyclic preferences can cause efficiency issues with the naive Borda Count method. To further refine the ranking, we devise a re-ranking stage, where the logits generated by the PPRM are used for soft comparison among nodes with equal Borda counts. Specifically, for two nodes $v_{i}$ and $v_{j}$ with equal Borda counts, the soft comparison rule can be denoted as $v_{i}\succ v_{j}\iff P(a_{i}\succ a_{j})>P(a_{j}\succ a_{i})$. This process ensures that the ranking remains consistent and reasonable even in the presence of cycles or local ambiguities.
@@
Global quantile score of solutions. Finally, the ranking is converted into a global quantile score $Q_{g}$ for each solution $v$ is $Q_{g}(v)=1-\frac{\text{rank}(v)-1}{|V|-1}$, where $\text{rank}(v)$ is the position of $v$ in the ranking based on Borda counts, and $|V|$ is the total number of nodes. To reflect local advantages in the search tree structure, the local win rate $Q_{l}(v)$ for a node $v$ is calculated in $\mathbf{C}$ with its children nodes $\text{Children}_{v}$ as follows:
@@
Finally, score $Q(v)$ for a solution is a weighted sum of local win rate $Q_{l}(v)$ and global quantile score $Q_{g}$.
Table 1: Performance comparison of models across benchmarks of different difficulties, as represented by GaoKao2023En (Liao et al., 2024), College Math (Tang et al., 2024), and OlympiadBench (He et al., 2024), which range from high school to Olympiad levels. Scores denoted with subscripted notations, such as maj@8, represent specific metrics, with major@8 as an example. Scores without subscripted notations reflect the model’s greedy performance evaluated in a zero-shot CoT manner.
Qwen2-7B-Instruct (Yang et al., 2024a) 85.7 52.9 36.4 21.3 24.5 68.2
Qwen2-72B-Instruct (Yang et al., 2024a) 93.2 69.0 58.7 33.2 43.2 84.4
DeepSeekMath-7B-RL (Shao et al., 2024) 88.2 52.4 43.6 19.0 37.5 64.8
Mathstral-7B-v0.1 (Mistral AI, 2024) 84.9 56.6 46.0 21.5 33.7 64.0
Qwen2-Math-7B-Instruct (Yang et al., 2024a) 89.9 75.1 62.1 38.2 45.9 63.8
Qwen2-Math-72B-Instruct (Yang et al., 2024a) 96.7 84.0 68.3 43.0 47.9 79.9
Meta-Llama-3.1-8B-Instruct (Meta, 2024b) + LLaMA-Berry (Ours)@8 $89.8_{\texttt{maj@8}}$ $54.8_{\texttt{maj@8}}$ $36.4_{\texttt{maj@8}}$ $24.8_{\texttt{maj@8}}$ $36.4_{\texttt{maj@8}}$ $68.3_{\texttt{maj@8}}$
$94.9_{\texttt{rm@8}}$ $69.4_{\texttt{rm@8}}$ $61.6_{\texttt{rm@8}}$ $47.2_{\texttt{rm@8}}$ $63.7_{\texttt{rm@8}}$ $82.9_{\texttt{rm@8}}$
+LLaMA-Berry (Ours)@16 $96.1_{\texttt{rm@16}}$ $75.3_{\texttt{rm@16}}$ $68.6_{\texttt{rm@16}}$ $55.1_{\texttt{rm@16}}$ $68.9_{\texttt{rm@16}}$ $88.3_{\texttt{rm@16}}$
@@ 3.1 Experiment Settings
Settings. To better evaluate the effectiveness of our approach, we select LLaMA-3.1-8B-Instruct model (Meta, 2024b) as the base model for SR-MCTS, without any additional training. We also train a Gemma2-2B-Instruct model (Google, 2024) as PPRM to provide reward signals during the search process. We develop the Berry-Tree inference framework to ensure robust and efficient inference, which supports advanced features, including fault tolerance, checkpoint recovery, multi-query concurrency, and automatic multi-server load balancing. Hyper-parameter settings are detailed in Appendix A.
@@
Grading. We evaluate algorithm-generated answers using the correctness evaluation standards as in Lightman et al. (2023), focusing on format adherence and content accuracy. The model is provided with a prompt specifying the expected answer format. We score answers as consistent if they exactly match the ground truth, closely approximate it numerically, or are equivalent in symbolic form. To ensure a comprehensive and rigorous evaluation, we adopt major@k (Kuncheva, 2014) and rm@k (Yang et al., 2024c), which can be unified as the solved rate of problems (Lightman et al., 2023; Luo et al., 2024a). The evaluation methods and metric details are further outlined in Appendix B.
3.2 Benchmarks
@@
General mathematical reasoning benchmarks. We summarize the results on general mathematical reasoning benchmarks in Table 1, which indicates that our method significantly boosts the base model’s performance. The results consistently demonstrate improvements across various levels of difficulty. Specifically, the solved rate of problems in 16 rollouts of Meta-Llama-3.1-8B-Instruct (Meta, 2024b) has been improved by more than 35% on three benchmarks. Qwen2-Math-72B-Instruct (Yang et al., 2024b) exhibits the strongest mathematical reasoning capability among the competing methods, while our LLaMA-Berry, built on a base model with only 8B parameters, exceeds it in the solved rate of problems on four benchmarks. In particular, LLaMA-Berry reaches 55.1% on OlympiadBench and 68.9% on College Math, surpassing it by 11.9% and 21%, respectively.
Cutting-edge mathematical Olympiad benchmarks. In Table 2, we compare the performance of LLaMA-Berry with other leading models on Olympic-level benchmarks. The results demonstrate that LLaMA-Berry is highly competitive on these benchmarks, demonstrating its capability in complex reasoning. Notably, on the most challenging AIME2024 benchmark, our method boosts the base model’s solving rate from 2/30 to 8/30, surpassing typical open-source models and commercial closed-source models, except for the OpenAI o1 series (OpenAI, 2024).
In addition to excelling in mathematical reasoning, our approach also excels across various science and engineering domains. For example, it achieves top performance on benchmarks such as MMLU STEM (Hendrycks et al., 2021a) in Table 1 and GPQA diamond (Rein et al., 2024) in Table 2. This demonstrates the method’s robustness and versatility, enabling it to tackle a wide range of technical challenges and highlighting its potential for broader applications in both research and practical scenarios.
@@
Table 2: Performance comparison across multiple olympiad benchmarks, including AIME24 (MAA, 2024), AMC23 (MAA, 2023), Math Odyssey (Fang et al., 2024), and GPQA Diamond (Rein et al., 2024).
@@
[…]
Claude 3 Opus 6.7 42.0 40.6 50.4
GPT 4 Turbo 3.3 47.0 38.8
GPT 4o 13.4 56.1
OpenAI o1 Preview 56.7 78.3
OpenAI o1 83.3 78.0
Gemini 1.5 Pro 6.7 45.0
Gemini Math-Specialized 1.5 Pro 23.3 55.8
Meta-LLaMA-3.1-8B-Instruct 6.7 15.7 41.7 30.4
Qwen2-Math-7B-Instruct 13.3 62.5
NuminaMath-72B CoT 3.3 52.5
Qwen2-Math-72B-Instruct 20.0 60.0
Meta-Llama-3.1-8B-Instruct + LLaMA-Berry (Ours)@8 $13.3_{\texttt{maj@8}}$ $22.9_{\texttt{maj@8}}$ $44.2_{\texttt{maj@8}}$ $39.4_{\texttt{maj@8}}$
$16.7_{\texttt{rm@8}}$ $48.2_{\texttt{rm@8}}$ $60.4_{\texttt{rm@8}}$ $77.3_{\texttt{rm@8}}$
+LLaMA-Berry (Ours)@16 $26.7_{\texttt{rm@16}}$ $54.2_{\texttt{rm@16}}$ $65.0_{\texttt{rm@16}}$ $92.4_{\texttt{rm@16}}$
Comparison with other tree-based or CoT methods. We compare our algorithm with other tree-based reasoning methods and CoT-based methods on GSM8K Cobbe et al. (2021), GSMHard (Gao et al., 2022), and MATH500 (Lightman et al., 2023) which is a representative and highly challenging 10% subset of MATH Hendrycks et al. (2021b) benchmark. As shown in Table 3, the performance of RAP Hao et al. (2023) and ToT (Yao et al., 2023) tends to degrade relative to more straightforward methods like Few-shot CoT and One-turn Self-Refine when the difficulty increases from GSM8K to GSMHard. We suspect the reason could be the weak self-evaluation capability of LLMs, which may guide reasoning steps to the inefficient side. Moreover, tree-based methods can incur more computational overhead than straightforward methods. In contrast, rStar Qi et al. (2024) and our method maintain a positive output performance trend, highlighting both approaches’ higher search efficiency.
To make a fair comparison between the reported results on LLaMA-3-8B-instruct from rStar Qi et al. (2024), self-consistency (Wang et al., 2022), and our algorithm, we also utilize the LLaMA-3-8B-instruct as the base model instead of the 3.1 version. We observe that our approach achieves on-par or even better performance with fewer rollouts. Specifically, our method achieves an accuracy of 88.1%, 31.5%, and 39.6% on GSM8K, GSMHARD, and MATH500 benchmarks, respectively, using the majority voting metric, which is among the same accuracy level as others while only consuming 1/2 of the rollout times of rStar and 1/8 of Self-consistency. This provides compelling validation of the efficacy of the EBC method and the aggressive exploration fostered by the dynamic pruning strategy.
Table 3: Performance of different tree-based methods for LLaMA-3-8B-Instruct on GSM8K, GSMHARD, and MATH500 benchmarks.
Method Benchmark GSM8K GSMHARD MATH500
@@
Self-Cons@8 78.4 28.5 30.0
Self-Cons@64 83.2 30.3 33.0
Self-Cons@128 84.7 31.2 33.8
ToT@32 69.1 19.6 13.6
RAP@32 80.6 29.6 18.8
rStar@32 $88.7_{\texttt{maj@32}}$ $33.4_{\texttt{maj@32}}$ $38.3_{\texttt{maj@32}}$
LLaMA-Berry (Ours)@8 $86.4_{\texttt{maj@8}}$ $30.2_{\texttt{maj@8}}$ $35.2_{\texttt{maj@8}}$
$94.1_{\texttt{rm@8}}$ $37.1_{\texttt{rm@8}}$ $56.4_{\texttt{rm@8}}$
LLaMA-Berry (Ours)@16 $88.1_{\texttt{maj@16}}$ $31.5_{\texttt{maj@16}}$ $39.6_{\texttt{maj@16}}$
$96.4_{\texttt{rm@16}}$ $41.1_{\texttt{rm@16}}$ $63.8_{\texttt{rm@16}}$